One mole of an ideal gas is first compressed isothermally at 380 K to half of its initial volume. Then 450 J of heat is added to it at constant volume. So, what is the network and total change in internal energy after 2 processes?
1
Expert's answer
2019-06-04T12:07:49-0400
  1. Isothermal compression:
"W_1=nRT \ln{\frac{V_f}{V_i}}"

"W_1=(1)(8.31)(380)\ln{\frac{1}{2}}=-2190\ J"

"\Delta U_1=0"

2.


"W_2=0"

"\Delta U_2=Q=450\ J"

Thus,


"W=W_1=-2190\ J"

"\Delta U=\Delta U_2=450\ J"